Andrea requires 30 minutes to swim 15 miles downstream. The return trip upstream takes 50 minutes.What is Andrea's speed in still water? How fast is the curret in miles per hour?
Answer by mananth(16946)

swimming speed =x mph
current speed =y mph
against current 0.75 hours
with current 0.5 hours

Distance with current 15 miles distance against current 15
t=d/r against current (x-y)
15.00 / ( x - y )= 0.75

0.75 x - -0.75 y = 15.00 ....................1
with current (x+y)
15.00 / ( x + y )= 0.50
0.50 ( x + y ) = 15.00
0.50 x + 0.50 y = 15.00 ...............2
Multiply (1) by 0.50
Multiply (2) by 0.75
we get 2.00
0.38 x + -0.38 y = 7.50
0.38 x + 0.38 y = 11.25
0.75 x = 18.75
/ 0.75
x = 25 mph

plug value of x in (1) y
0.75 x -0.75 y = 15
18.75 -0.75 -18.75 = 15
-0.75 y = 15
-0.75 y = -3.75 mph
y = 5
swimming 25 mph
current 5 mph